On a train, it took 3/4 hour to travel a distance of 52 1/2 miles. What was the average speed of the train?
Answer by nyc_function(2733)

r = D/t

Solve for r when D = 52_1/2 and t = 3/4



We can write 52_1/2 as 105/2.

r = 105/2 divided by 3/4

Can you finish?